OpenStudy (anonymous):

3x + 7 = 22 what's x? 3x + 7 = 22 what's x? @Mathematics

OpenStudy (anonymous):

subtract 7 to both sides: 3x+7=22 -7 -7 ------------- 3x=15 divded by three on each side 3x=15 x=5 -- --- 3 3